OpenStudy (anonymous):

Piper invested $7,400 in a Certificate of Deposit, $4,980 in a corporate bond, $5,100 in a common stock, and $6,350 in a preferred stock, The COD has a rate of return of 3.4%; the corporate bond is 5.8%; the common stock’s rate of return is –1.9%; and the preferred stock has a rate of return of 4.6%. What is Piper’s weighted mean rate of return? 4.60% 3.90% 2.98% 3.09%

OpenStudy (tkhunny):

\(\dfrac{Sum\;of\;all\;returns}{Sum\;of\;all\;imvestments} = ??\)

OpenStudy (anonymous):

\[$7,400 + $4,980 + $5,100 + $6,350 / 3.4% + 5.8% + -1.9% + 4.6%\]

OpenStudy (tkhunny):

?? What is that? What is the return AMOUNT on the COD?

OpenStudy (anonymous):

0.034

OpenStudy (tkhunny):

No, that is the return RATE. You need the return AMOUNT. Give it another go.

OpenStudy (tkhunny):

7400*0.034 = ??

OpenStudy (anonymous):

251.6

OpenStudy (tkhunny):

Okay, now the other three.

OpenStudy (anonymous):

The other one I got is 288.84

OpenStudy (tkhunny):

The other ONE? Please read and understand the problem statement. No mope guessing. Just move along one piece at a time. COD: $7,400 * 0.034 = CorpBond $4,980 * 0.058 = CommStock: $5,100 * (-0.019) = PrefStock: $6,350 * 0.046 = Calculate ALL FOUR of those and you will have the total investment income. Yes, the Common Stock will be negative. The calculate: \(\dfrac{Total\;Investment\;Return}{Total\;Investment}\)